# Endmember Extraction from Hyperspectral Images Using Self-Dictionary Approach with Linear Programming

### v2

@@ Appendix B Cluster Construction in Step 2-1 of Algorithm 1
Let $\mathcal{S}\in\Omega_{i}$ and $\bm{p}\in\mathbb{R}^{n}_{+}$. We regard $\bm{p}$ as a point list for $\mathcal{S}$ and define the score of $\mathcal{S}$ based on the point list $\bm{p}$ as follows:
@@
Step 2-1 of Algorithm 1 constructs $\mathcal{S}_{\ell}$ as follows:
@@
Appendix C Proof of Theorem 1
Proof:
Let $(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{M})$. Then, $X^{*}$ is a feasible solution of $\mathsf{H}(\mathcal{L},\mathcal{M})$. Thus, we get $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))\leq\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}$. Furthermore, taking into account that $(X^{*},F^{*},G^{*},u^{*})$ satisfies all constraints of $\mathsf{P}(\mathcal{L},\mathcal{M})$, we find that
$\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}$ $=\|F^{*}-G^{*}\|_{1}$
$=\max_{j=1,\ldots,m}\|F^{*}(:,j)-G^{*}(:,j)\|_{1}$
$\leq\max_{j=1,\ldots,m}\|F^{*}(:,j)\|_{1}+\|G^{*}(:,j)\|_{1}$
$=\max_{j=1,\ldots,m}\sum_{i=1}^{d}F^{*}(i,j)+G^{*}(i,j)$
$\leq u^{*}=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M})).$
Accordingly,
Let $\hat{X}^{*}$ be the optimal solution of $\mathsf{H}(\mathcal{L},\mathcal{M})$, and let us use it to construct $R=A(\mathcal{M})\Pi-A(\mathcal{L})\hat{X}^{*}$. Note that $R$ satisfies the relation $\|R\|_{1}=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$, which will be used below. We show that $(\hat{X}^{*},R^{+},R^{-},\|R\|_{1})$ is a feasible solution of $\mathsf{P}(\mathcal{L},\mathcal{M})$. Obviously, it satisfies all the constraints except the second one. Here, recall equality (1) shown in Section II-A For the second constraint, this equality implies that
$\sum_{i=1}^{d}R^{+}(i,j)+R^{-}(i,j)\leq\|R\|_{1}.$
Accordingly,
$\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))=\|R\|_{1}\geq\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M})).$ (14)
In addition, as explained just before this theorem, the duality theorem ensures
$\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M})).$ (15)
From (13), (14) and (15), we obtain the equalities,
$\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$
$=\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}=\|R\|_{1}.$
As shown above, $X^{*}$ and $(\hat{X}^{*},R^{+},R^{-},\|R\|_{1})$ are feasible solutions of $\mathsf{H}(\mathcal{L},\mathcal{M})$ and $\mathsf{P}(\mathcal{L},\mathcal{M})$, respectively. Hence, this completes the proof of parts (i)-(iii). ∎
Appendix D Proof of Lemma 1
Proof:
We plug in the optimal solution of $\mathsf{D}(\mathcal{L},\mathcal{L})$ for the first constraint and then express the resulting inequality as $M\leq O$ by letting
The trace of $M$ is given as
by $M\leq O$ and $Z^{*}\geq O$. We also have $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))\geq 0$. Indeed, $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by Theorem 1 (i), and $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))\geq 0$, since the objective function value of $\mathsf{H}(\mathcal{L},\mathcal{L})$ always takes a nonnegative value. We thus get
Appendix E Choosing the Value of $\omega$
Our data-specific preprocessing for the Urban HSI dataset requires us to choose two parameters $\phi$ and $\omega$. In particular, as shown in Figure 4, the endmember extraction performance of the methods depends on the choice of $\omega$. When drawing the histogram of density $\rho$ versus the number of pixels while varying $\phi$ from $0.45$ to $0.6$ in $0.05$ increments, as in the case where $\phi=0.4$, a small peak appeared in a low-density range. Therefore, on the basis of the experimental results for $\phi=0.4$, we can describe the criteria for choosing $\omega$ when $\phi$ is between $0.45$ and $0.6$ as follows.
@@ Appendix E Choosing the Value of $\omega$
We chose $\omega$ by following criteria (1) and (2) and ran EEHT-C and six existing methods on Urban. Table III summarizes the average MRSA values. We see that the average MRSA values of EEHT-C are nearly unchanged for $\phi$ from $0.4$ to $0.6$. These results show that criteria (1) and (2) for choosing $\omega$ can yield consistent endmember extraction results for EEHT-C on Urban when $\phi$ belongs to this range.
TABLE III: Average MRSA values (×$\times 10^{2}$) of methods run on Urban while varying $\phi$ and using criteria (1) and (2) for the choice of $\omega$. The best scores are highlighted in bold.
$\phi$ $\omega$ EEHT-C MERIT SPA PSPA ER VCA SNPA
0.40 0.10 7.9 19.5 17.8 13.0 17.0 28.0 19.7
0.45 0.15 8.4 17.0 17.8 17.0 17.0 28.3 19.7
0.50 0.30 9.3 17.0 17.8 17.0 17.0 28.3 19.7
0.55 0.45 8.2 11.9 17.8 17.0 17.0 28.0 19.7
0.60 0.60 8.2 17.0 17.8 17.0 17.0 28.3 19.7
Appendix F Figures
Figure 5 displays the RGB images of three datasets: Jasper Ridge, Samson, and Urban. Figure 6 presents the ground truth and abundance maps obtained by EEHT-C and six existing methods for the Urban dataset with $\phi=0.4$ and $\omega=0.1$. The ground truth was derived by using the reference signatures.
Figure 5: RGB images of Jasper Ridge (top-left), Samson (top-right), and Urban (bottom).

### v3

@@ Appendix D Cluster Construction in Step 2-1 of Algorithm 1
Let $\mathcal{S}\in\Omega_{i}$ and $\bm{p}\in\mathbb{R}^{n}_{+}$. We regard $\bm{p}$ as a point list for $\mathcal{S}$ and define the score of $\mathcal{S}$ with respect to the point list $\bm{p}$ as
@@
Step 2-1 of Algorithm 1 then constructs $\mathcal{S}_{\ell}$ as follows:
@@
Appendix E Effect of Parameters $\zeta$ and $\eta$ on the Computational Time of RCE-SR
We conducted experiments to investigate the effect of the parameters $\zeta$ and $\eta$ on the computational time of RCE-SR. In these experiments, $\zeta$ was varied from $5$ to $20$ in increments of $5$, and $\eta$ was varied from $50$ to $200$ in increments of $50$, resulting in $16$ combinations of $(\zeta,\eta)$. For each combination of $(\zeta,\eta)$, RCE-SR was applied to six datasets, comprising the five datasets used in the previous experiments and an additional dataset with $n=5000$. For each dataset, the elapsed time of RCE-SR for $(\zeta,\eta)=(10,100)$ was used as the baseline and set to $1$, and the times for the other combinations were evaluated relative to this baseline. Table VII summarizes the results, showing that even when $(\zeta,\eta)$ were varied across the $16$ combinations, the elapsed time of RCE-SR remained within the range of $0.8$ to $1.8$.
TABLE VII: Minimum and maximum elapsed times of RCE-SR, normalized by the baseline $(\zeta,\eta)=(10,100)$, for the $16$ combinations of $(\zeta,\eta)$.
$n=500$ $n=1000$ $n=1500$ $n=2000$ $n=2500$ $n=5000$
Min 0.96 0.84 0.82 0.92 0.98 1.01
Max 1.72 1.40 1.38 1.48 1.62 1.22
TABLE VIII: MRSA scores ($\times 10^{2}$) on Urban, obtained by varying $\phi$ and choosing $\omega$ according to criteria (1) and (2). The best scores are highlighted in bold.
$(\phi,\omega)$ EEHT-C MERIT SPA PSPA ER VCA SNPA
(0.40, 0.10) 7.9 19.5 17.8 13.0 17.0 28.0 19.7
(0.45, 0.15) 8.4 17.0 17.8 17.0 17.0 28.3 19.7
(0.50, 0.30) 9.3 17.0 17.8 17.0 17.0 28.3 19.7
(0.55, 0.45) 8.2 11.9 17.8 17.0 17.0 28.0 19.7
(0.60, 0.60) 8.2 17.0 17.8 17.0 17.0 28.3 19.7
Appendix F Effect of Parameter $\omega$ on Endmember Extraction Performance
Our data-specific preprocessing for the Urban HSI dataset requires us to choose two parameters $\phi$ and $\omega$. In particular, as shown in Figure 7, the endmember extraction performance of the methods depends on the choice of $\omega$. When plotting the histogram of density $\rho$ versus the number of pixels while varying $\phi$ from $0.45$ to $0.6$ in increments of $0.05$, as in the case where $\phi=0.4$, a small peak appeared in a low-density range. Therefore, on the basis of the experimental results for $\phi=0.4$, we can describe the criteria for choosing $\omega$ when $\phi$ is between $0.45$ and $0.6$ as follows.
@@ Appendix F Effect of Parameter $\omega$ on Endmember Extraction Performance
We chose $\omega$ by following criteria (1) and (2) and ran EEHT-C and six existing methods on Urban. Table VIII summarizes the MRSA scores. We see that the MRSA scores of EEHT-C are nearly unchanged for $\phi$ from $0.4$ to $0.6$. These results demonstrate that criteria (1) and (2) for choosing $\omega$ yield consistent endmember extraction results for EEHT-C on Urban when $\phi$ belongs to this range.